5. Add 21+27 using 6-bit signed binary numbers.
Answer #1

Magnitude of 21 in binary= 10101
Magnitude of 27 in binary= 11011

Now using 6-bit signed binary representation additiona we get :
            11111
+21 = 010101
+27 = 011011
-------------------------
            110000

(110000)2 = 48

So 21+27=48 ;

Note:-

It's an invalid operation(using 6bit signed binar y operation) as we can only represent -32 to +31 using 6 bit signed binary representation. But here the sum is 48 out of the range of 6 bit signed binary representation.